# Explainability of Machine Learning Approaches in Forensic Linguistics: A Case Study in Geolinguistic Authorship Profiling

### v1

@@ Abstract
Forensic authorship profiling uses linguistic markers to infer characteristics about an author of a text. This task is paralleled in dialect classification, where a prediction is made about the linguistic variety of a text based on the text itself. While there have been significant advances in the last years in variety classification (Jauhiainen et al., 2019) and state-of-the-art approaches reach accuracies of up to 100% depending on the similarity of varieties and the scope of prediction (e.g., Milne et al., 2012; Blodgett et al., 2017), forensic linguistics rarely relies on these approaches due to their lack of transparency (see Nini, 2023), amongst other reasons. In this paper we therefore explore explainability of machine learning approaches considering the forensic context. We focus on variety classification as a means of geolinguistic profiling of unknown texts. For this we work with an approach proposed by Xie et al. (2024) to extract the lexical items most relevant to the variety classifications. We find that the extracted lexical features are indeed representative of their respective varieties and note that the trained models also rely on place names for classifications.
@@ 1 Introduction
Forensic authorship analysis, i.e., the analysis of texts to gain information about the author(s) of those texts, is a central domain of research within forensic linguistics. Generally, there are three main areas within authorship analysis working to infer information from questioned documents: Firstly, authorship attribution is interested in learning who the most likely candidate author is given a comparison with the writings of a set of potential authors; Secondly, authorship verification focuses on evaluating if texts were authored by the same person; And thirdly, authorship profiling works to infer characteristics of an author by their language use (Roemling and Grieve, 2024).
While authorship attribution and verification rely on comparisons between (groups of) texts, profiling works only with the questioned document(s) itself. Generally, all areas of authorship analysis can be approached qualitatively and/or quantitatively if the amount of data available permits quantitative analysis. For example, quantitative work in authorship profiling has researched inferring age or gender (Nini, 2018) or native language influence (Perkins and Grant, 2018) from questioned documents.
Nevertheless, forensic authorship profiling is often done in a manual or qualitative way, relying on the expertise of the forensic linguist as can be seen in famous examples like the Unabomber analysis (Leonard et al., 2017) or the devil strip ransom note (Shuy, 2001). In both cases, law enforcement was interested in the regional origin as one of the relevant characteristics of the authors. This background can be inferred through analysing the regional linguistic variation, i.e., the use of regional dialect, in the questioned documents. This is referred to as regional or geolinguistic profiling (Roemling and Grieve, 2024) and is a parallel task to inferring the regional variety of a document as is done in dialect classification or language variety identification (Jauhiainen et al., 2019).
Even though research in forensic linguistics works more and more with statistical and computational approaches (e.g., Bevendorff et al., 2023; Ishihara et al., 2024), authorship profiling often remains a manual task. This is at times credited to the black-box approaches in current NLP research, meaning that the lack of explainability precludes these approaches from being used in legal settings (see Nini, 2023).
@@ 2 Related Work
Xie et al. (2024) were the first to research explainability in a dialect classification context. In order to analyze the classifier, they extracted lexical features that were highly relevant to the classification aiming to use this knowledge for dialect research and not only as a means to understand machine learning approaches. They stated that they relied on lexical items, owing to the complex nature of handling preprocessing like parsing or POS-tagging in low(er)-resourced varieties and indicated that refined approaches would be beneficial. However, previous research on regional variation using social media data has shown that approaches using lexical features provided excellent results (e.g., Doyle, 2014; Huang et al., 2016; Eisenstein, 2017; Grieve et al., 2018, 2019).
@@
In a forensic authorship profiling setting, an approach like this seems beneficial. It reaches high accuracies in language variety identification, while similarly providing explanations of which features predominantly influenced the classification. The approach also has the advantage of eliminating or at least minimising researcher bias as the models learn the relevant features themselves, whereas it is the forensic linguist’s expertise that culls the features in a qualitative analysis (see Grant and Grieve, 2022) [^1] Note, however, that the training data itself may introduce bias into the model (see Blodgett et al., 2020).. While the approach does not fully explain the inner workings of the model, experts can use the extracted features to a) verify that the model indeed reached a sound decision, for example by evaluating the features against previous dialectological findings, and b) use the explanations to introduce the method to law enforcement or jurisprudence. Even if the classifiers themselves do not meet court admissibility standards (Coulthard, 2013; Hammel, 2022), extracted features can be used for authorship work to contribute to a more objective analysis.
@@ 3 Data
We work with a corpus of German social media data, which was collected by Hovy and Purschke (2018); Purschke and Hovy (2019) from the platform Jodel and which has also been used in VarDial classification tasks (Gaman et al., 2020; Chakravarthi et al., 2021). Jodel is structurally similar to Twitter/X, however it only allows anonymous posts. Users of Jodel can interact with other users in a 10-15 km radius around their own location, so all posts are geolocated. The corpus contains posts from Austria, Germany and Switzerland. While most of the data is written in standard German, it shows clear regional patterns. Especially in Switzerland, Austria and, in parts, Bavaria writing is considerably further from standard German (Purschke and Hovy, 2019). Posts from Romandy contain substantial amounts of French.
The corpus consists of approximately 240 million tokens from about 8500 locations, however only 388 locations have a token count of over 10k. For our classification experiments, we mapped these locations onto wider dialect regions following three settings with 3, 4 and 5 classes respectively. The 3-class distinction is based on national borders, so the classes reflect Austria, Germany and Switzerland. In the 4-class setting Germany is additionally divided into two parts, north and south (at latitude 50.33° N) and for the 5-class setting the southern region of Germany is further split into east and west (at longitude 9.97° E). These divisions were operationalized based on knowledge from traditional dialectology (Wiesinger, 1983; König, 2004).
We randomly sampled 200k examples per class for training, and 20k examples per class for development and testing, respectively. Besides some simple whitespace normalization (i.e., removing line breaks and tabs inside a message), we did not perform any preprocessing.
@@ 4.2 Leave-one-word-out classification
The steps described above produce explanations at instance level, i.e., the most impactful words of each instance. This can be an interesting result in itself, but for the analysis reported here, we aggregate the explanations across all instances of the test set. The resulting list is then processed in the following way:
@@
    For each class, we select the 100 words with the highest impact scores for the analysis.
@@ 5 Results
Following the LOO method, we produce a list of 100 words with the highest impact scores per class in all settings. A manual inspection of the lists quickly shows a prevalence of place names and related items like Zürich or Österreicher (G ‘Austrian’). Therefore, as a first step, we count the amount of these words among the top 100 and find that, on average, 14% of words are local references. In terms of classification, these results are expected (see, e.g., Nasar et al., 2022) and it is apparent how these words are indicative of location given what they denote. Although it does not take a forensic linguist to understand the connection of these items to location, it is noteworthy that the classification models pick up on these words and that they match the region they are impactful for.
@@
Given that the results are similar for classes based on country borders across the three settings, we focus our analysis on the 5-class setting. As a reminder, in this setting, Austria and Switzerland form individual classes, whereas Germany is split into three regions (northern, south-eastern and south-western). Generally, the results show that a large proportion of extracted words are regionalized and some are prototypical dialect items.
@@
For Switzerland, we find both Swiss German and French lexical items to be the most impactful, such as Kei (SWG ‘not a/no’), Isch (SWG ‘is’) and bim (SWG ‘at the’), or pourquoi (FR ‘why’), Avec (FR ‘with’) and raison (FR ‘reason’). Whereas for Austria, the data shows that some of the most impactful lexical items are Oasch (AG ‘ass’), Gspusis (AG ‘affairs’), and Matura (AG ‘high-school (diploma)’). These three items are examples of a textualization of regional pronunciation, a regional item and a standard Austrian variant, respectively. Also, items that are relevant to the Jodel demographic, such as Jus (AG ‘law (studies)’), are extracted. This indicates that an automated classification between the three countries, for instance to clarify jurisdiction, seems reasonable.
@@
For the German classes we find that a large proportion of top words appear textualized in standard German as opposed to more colloquial spellings including abbreviations and ellipses, which we may expect. Examples of this include Dankeschön (G ‘thank you’) or Vorname (G ‘first name’). Considering the division within Germany, we find that several forms of the verb gucken (G ‘to look/watch’) are impactful for the northern class, which is a variant we know to be regionalized and appearing in varieties in central and northern Germany (König, 2004, p. 235). For the south-east class the data shows items like Ritter (G ‘knight’), #traudel (female first name) and local beer types. For the south-west we find that items identified as relevant by Purschke and Hovy (2019), are also extracted by the LOO model, like Möppes (G ‘breasts’ or ‘female user’) and Lörres (G ‘penis’ or ‘male user’), although the authors argue that these forms are more Jodel- than region-specific. These items are also impactful for the Germany-class in the 3-class setting.
@@ Acknowledgements
Dana Roemling was supported by the UKRI ESRC Midlands Graduate School Doctoral Training Partnership ES/P000711/1.
We thank Dirk Hovy and Christoph Purschke for sharing their data with us.
We wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.

### v2

@@ Abstract
Forensic authorship profiling uses linguistic markers to infer characteristics about an author of a text. This task is paralleled in dialect classification, where a prediction is made about the linguistic variety of a text based on the text itself. While there have been significant advances in recent years in variety classification, forensic linguistics rarely relies on these approaches due to their lack of transparency, among other reasons. In this paper we therefore explore the explainability of machine learning approaches considering the forensic context. We focus on variety classification as a means of geolinguistic profiling of unknown texts based on social media data from the German-speaking area. For this, we identify the lexical items that are the most impactful for the variety classification. We find that the extracted lexical features are indeed representative of their respective varieties and note that the trained models also rely on place names for classifications.
@@ 1 Introduction
Forensic authorship analysis is a key area of research within forensic linguistics that seeks to gain information about the author(s) of a text. Generally, there are two central domains of research within authorship analysis: comparative authorship analysis uses various methodologies to compare questioned and known documents to attribute authorship, while sociolinguistic or authorship profiling relies on the analysis of questioned texts alone to infer characteristics of an author (Grant, 2022; Roemling and Grieve, 2024). Both areas of authorship analysis can be approached qualitatively and/or quantitatively if the amount of available data permits it. For example, quantitative work in authorship profiling has researched inferring age or gender (Nini, 2018) or native language influence (Kredens et al., 2019) from questioned documents.
Nevertheless, forensic authorship profiling is often carried out in a manual or qualitative way, relying on the expertise of the forensic linguist. This is evident in famous examples like the Unabomber analysis (Leonard et al., 2017) or the devil strip ransom note (Shuy, 2001). In both cases, law enforcement was interested in the regional origin of the authors. This background can be inferred through analyzing the regional linguistic variation, i.e., the use of regional dialect, in the questioned documents. This is referred to as regional or geolinguistic profiling (Roemling and Grieve, 2024) and is a task parallel to inferring the regional variety of a document as is done in language identification (Jauhiainen et al., 2019).
Even though research in forensic linguistics works more and more with statistical and computational approaches (e.g., Bevendorff et al., 2023; Ishihara et al., 2024; Nini et al., 2024), authorship profiling often remains a manual task. This is at times credited to the black-box approaches in current NLP research, meaning that the lack of explainability precludes these approaches from being used in legal settings (see Nini, 2023).
@@ 2 Related work
Xie et al. (2024) were the first to research explainability in a dialect classification context. In order to analyze the classifier, they extracted lexical features that were highly relevant to the classification, aiming to use this knowledge for dialect research and not only as a means to explore machine learning approaches. They relied on lexical items, owing to the complex nature of handling preprocessing like tokenisation or POS-tagging in low(er)-resourced varieties. They indicated that refined approaches would be beneficial. However, previous research on regional variation using social media data has shown that approaches using lexical features provided excellent results (e.g., Doyle, 2014; Huang et al., 2016; Eisenstein, 2017; Grieve et al., 2018, 2019).
@@
In a forensic authorship profiling setting, an approach like this appears beneficial as it reaches high accuracies in language identification, while similarly providing explanations by extracting the features that influenced the classification. While the original study focused on improving research methods in dialectology, we apply the method to evaluate its usefulness in a forensic context. Additionally, the approach has the advantage of eliminating or at least minimising researcher bias as the models learn the relevant features themselves, whereas it is the forensic linguist’s expertise that culls the features in a qualitative analysis (see Grant and Grieve, 2022) [^1] Note, however, that the training data itself may introduce bias into the model (see Blodgett et al., 2020).. While the approach does not fully explain the inner workings of the model, experts can use the extracted features to a) verify that the model indeed reached a sound decision, for example by evaluating the features against previous dialectological findings, and b) use the explanations to introduce the method to law enforcement or jurisprudence. Even if the classifiers themselves do not meet court admissibility standards (Coulthard, 2013; Hammel, 2022), extracted features can be used for authorship work to contribute to a more objective analysis.
@@ 3 Data
We work with a corpus of German social media data from the platform Jodel. The corpus was collected by Hovy and Purschke (2018); Purschke and Hovy (2019). It has also been used in VarDial classification tasks (Gaman et al., 2020; Chakravarthi et al., 2021). Jodel is structurally similar to Twitter/X, however it only allows anonymous posts. Users of Jodel can interact with other users in a 10-15 km radius around their own location, so all posts are geolocated. The corpus contains posts from Austria, Germany and Switzerland. While most of the data is written in standard German, it shows clear regional patterns. Especially in Switzerland, Austria and, in parts, Bavaria writing is considerably further from standard German (Purschke and Hovy, 2019). Posts from Romandy contain substantial amounts of French. This data differs from the corpus used in the original study (see Xie et al., 2024) in terms of register and genre.
Figure 1: Operationalization of dialect regions
The corpus consists of approximately 240 million tokens from about 8500 locations, however only 388 locations have a token count of over 10k. For our classification experiments, we mapped these locations onto wider dialect regions following three settings with 3, 4 and 5 classes respectively. The 3-class distinction is based on national borders, so the classes reflect Austria, Germany and Switzerland. In the 4-class setting Germany is additionally divided into two parts, north and south (at latitude 50.33° N) and for the 5-class setting the southern region of Germany is further split into east and west (at longitude 9.97° E) (see Figure 1). These divisions were operationalized based on knowledge from traditional dialectology (Wiesinger, 1983; König, 2004).
We randomly sampled 200k posts per class for training, and 20k posts per class for development and testing, respectively. On average, a post contains 11.5 tokens. Besides some simple whitespace normalization (i.e., removing line breaks and tabs inside a post), we did not perform any preprocessing.
@@ 4 Methods and experiments
The basis for our analysis is the post-hoc LOO approach proposed by Xie et al. (2024). While we replicate most of the methodology, we mark any changes from the original in our explanations below. Crucially, and in contrast to the original dialectal research interest, we apply the approach with the goal of evaluating its usefulness in a forensic setting. Additionally, we work with data that is different in terms of register and genre and thus adds to our overall understanding of the approach.
@@ 4.2 Leave-one-word-out classification
The steps described above produce explanations at instance level, i.e., the most impactful words of each instance. In a forensic case setting with limited data, this could already be leveraged by using the impactful features in a qualitative analysis or simply evaluating and explaining the prediction made by the classifier. Consequently, this can be an interesting result in itself, but for the analysis reported here, we aggregate the explanations across all instances of the test set for evaluation. The resulting list is processed in the following way:
@@
    For each class, we select the 100 words with the highest average impact scores for the analysis.
@@ 5 Results
Following the method described above, we produce a list of 100 words with the highest impact scores per class in all settings. A manual inspection of the lists quickly shows a prevalence of place names and related items like Zürich or Österreicher (G ‘Austrian’) [^10] Throughout the remainder of the paper, the examples from Germany are marked with G, those from Austria with AG, and those from Switzerland with SWG. Examples from Romandy are marked FR for French.. Therefore, as a first step, we count the amount of these words among the top 100 and find that, on average, 14% of words are local references. In terms of classification, these results are expected (see, e.g., Nasar et al., 2022) and it is apparent how these words are indicative of location given what they denote. Although it does not take a forensic linguist to understand the connection of these items to location, it is noteworthy that the classification models pick up on these words and that they match the region they are impactful for.
@@
Given that the results are similar for classes based on country borders across the three settings, we focus our analysis on the 5-class setting. As a reminder, in this setting, Austria and Switzerland form individual classes, whereas Germany is split into three regions (see Figure 1). Generally, the results show that a large proportion of extracted words are regionalized and some are prototypical dialect items.
@@
For Switzerland, we find both Swiss German and French lexical items to be the most impactful, such as Kei (SWG ‘not a/no’), Isch (SWG ‘is’) and bim (SWG ‘at the’), or pourquoi (FR ‘why’), Avec (FR ‘with’) and raison (FR ‘reason’). Whereas for Austria, the data shows that some of the most impactful lexical items are Oasch (AG ‘ass’), Gspusis (AG ‘affairs’), and Matura (AG ‘high-school diploma’). These three items are examples of a textualization of regional pronunciation, a regional item and a standard Austrian variant, respectively. Also, items such as Jus (AG ‘law (studies)’), which are relevant to the Jodel demographic of mostly students under the age of 27, are extracted. This indicates that an automated classification between the three countries, for instance to clarify jurisdiction, seems reasonable.
@@
For the three German classes we find that a large proportion of top words appear textualized in standard German as opposed to more colloquial spellings including abbreviations and ellipses, which we may expect. Examples of this include Dankeschön (G ‘thank you’) or Vorname (G ‘first name’). Considering the division within Germany, we find that several forms of the verb gucken (G ‘to look/watch’) are impactful for the northern class, which is a variant we know to be regionalized and appearing in varieties in central and northern Germany (König, 2004, p. 235). For the south-east class the data shows items like Ritter (G ‘knight’), #traudel (female first name) and local beer types. For the south-west we find that items identified as relevant by Purschke and Hovy (2019), are also extracted by the LOO model, like Möppes (G ‘breasts’ or ‘female user’) and Lörres (G ‘penis’ or ‘male user’), although the authors argue that these forms are more Jodel- than region-specific. These items are also impactful for the Germany-class in the 3-class setting.
@@ Acknowledgements
Dana Roemling was supported by the UKRI ESRC Midlands Graduate School Doctoral Training Partnership ES/P000711/1. Yves Scherrer and Aleksandra Miletić were supported by the Academy of Finland through project No. 342859 “CorCoDial – Corpus-based computational dialectology”.
We thank Dirk Hovy and Christoph Purschke for sharing their data with us. We also wish to acknowledge CSC – IT Center for Science, Finland, for computational resources.